10

Village Finance Co. advanced three loans to Kamiko-$4,200 on June 20, $3,800 on August 25,and $3,300 on October 29, Simple interest at 8.25% was charged on all three loans, and all were repaid on December 31 when some bonds that she owned matured What total amount was required to pay off the loans? (Use 365 days a year. Do not round intermediate calculations and round your

final answer to 2 decimal places.)
1
Expert's answer
2021-09-20T23:25:23-0400

Loan 1 is 4200$ taken on June 20


Loan 2 is 3800$ taken on August 25


Loan 3 is 3300$ taken on October 29


Loans will be repaid on 31st December.

the interest will be charged till 30th December.

For loan 1 , the repayment of loan will be = June (11 days) including the 20th of June + July (31 days ) + Aug (31days) + Sep (30days) + Oct(31days) + Nov(30days) + Dec(30days)

= 194 days

So time =t= 194/365

Interest charged =r= 8.25% annually , that is for 365 days

Interest charged on loan for its repayment = P×r×t /100 =

(4200×8.25×194)/(365×100) = 184.17

Amount required to repay loan1 = 4200+184.17 = 4384.17 $


For loan2 = number of days = 7(Aug) + 30(Sep) + 31 (Oct) + 30 (Nov) + 30 (Dec) = 128 days

Time = 128/365

Interest =( 3800×8.25×128)/(100×365) = 109.94$

Amount required to repay loan2 = 3800+108.94= 3908.94


Similarly for loan3 repayment of loan = 3(Oct) + 30 (Nov) + 30 (dec) = 63 days

So interest charged = (3300×8.25×63)/(100×365) = 46.99

So amount required to repay loan3 = 3300+46.99 = 3346.99


Total amount required to payoff the loans = 4384.17+ 3908.94 + 3346.99 = 11640.1 $